OpenStudy (anonymous):

square root of -9 all over ( 3-2)i + (1+5i)

OpenStudy (anonymous):

\[\bf \frac{ \sqrt{-9} }{ 3-2i+1+5i }=\frac{ \sqrt{-1} \sqrt{9} }{3-2i+1+5i }\]Note that complex numbers follow all the same laws with addition and other operations like real numbers hence you can add like terms in the denominators just the way you would with real numbers have 'i' been a real variable. Also re-call \(\bf \sqrt{-1}=i\). Now I'll leave the simplifying to you :D @NickiSilva17

OpenStudy (anonymous):

@NickiSilva17 Can you simplify or do you require further help?

OpenStudy (anonymous):

Futher help please just started learning this so im practicing

OpenStudy (anonymous):

and thanks for your help

OpenStudy (anonymous):

more help pls

OpenStudy (usukidoll):

combine like terms on the bottom of the denominator

OpenStudy (anonymous):

is the answer 12-9i over 7?

OpenStudy (usukidoll):

 I got this